OpenStudy (anonymous):

identify the center and radius of the circle with the equation (x+5)^2+(y-2)^2=9 A. This equation represents a circle with center (5,-2) and a radius of 3 units B. This equation represents a circle with center (-5,2) and a radius of 3 units C. this equation represents a circle with center (5,-2) and a radius of 9 units D. This equation represents a circle with center (-5,2) and a radius of 9 units E. None

hartnn (hartnn):

equation of circle with center (h,k) and radius 'r' is \(\Large (x-h)^2 + (y-k)^2=r^2\)

hartnn (hartnn):

so compare (x+5)^2+(y-2)^2=9 with the above equation to get h,k,r

OpenStudy (anonymous):

would that be ?

OpenStudy (anonymous):

abcd or e?

hartnn (hartnn):

compare \(\Large x-h\) with \(\Large x+5\) can you tell me what will be the value of 'h' be ?

OpenStudy (anonymous):

told you @hartnn . he asked would it be ,a b c d or e

OpenStudy (anonymous):

helll help you @adugger935 (:

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

would h be 2?

hartnn (hartnn):

 try again ?